OpenStudy (anonymous):

Solve the system of equations. 2x + 7y = -38 -5x - 8y = 38 (-6, 2) (2, -6) {}

OpenStudy (anonymous):

2x + 7y = -38 ----------(1 -5x - 8y = 38 -----------(2 Multiplying eq (1 by 5 and adding them (2 by 2 we find 10x + 35y = -190 ----------(1 -10x - 16y = 76 -----------(2 ----------------------------------- 19y = - 144 y=-6 Substituting y=-6 in eq (1 we find 2x + 7*(-6) = -38 2x-42=-38 2x= 42-38 2x= 4 x= 2 Hence (2, -6) is the solution of the given system of equations.

OpenStudy (anonymous):

thanks it allmakes sense now